A nurse has a solution that is 25% alcohol and another that is 50% alcohol. How much of each must he use to make 45 liters of a solution that is 40% alcohol?
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A nurse has a solution that is 25% alcohol and another that is 50% alcohol. How much of each must he use to make 45 liters of a solution that is 40% alcohol?
------
Equation:
alch + alch = alch
0.25x + 0.50(45-x) = 0.40*45
----------
25x + 50*45 - 50x = 40*45
-----
-25x = -10*45
-----
x = 18 liters (amt. of 25% solution needed)
45-x = 27 liters (amt. of 50% solution needed)
==============================

Answer by Alan3354(69443)

A nurse has a solution that is 25% alcohol and another that is 50% alcohol. How much of each must he use to make 45 liters of a solution that is 40% alcohol?
------------
t = amount of 25%
f = amount of 50%
-----
t + f = 45
25t + 50f = 45*40